Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 561
Entry time: 12 Sep 2018, 14:52
|
Author: |
Adriano Violante |
Location: |
General |
Severity: |
Information |
Subject: |
Status report after dispatch meeting 09.30 |
|
|
SPB
- Testing of the 4 images preview mode for AGIPD will be done today.
- Interlock definitions are being updated. Loop update will be done after the users experiments.
FXE
SQS
- Problem with the digitizers bandwidth has been addressed by AE. ITDM has to reconfigure the DAQ. ==> Some issues still present. It seems because of lack of communication. AE/ITDM/CAS/SQS meeting tomorrow to clarify.
- Loop 6 being prepared. A couple of issues with Smaract and Technosoft controllers have to be addressed with the involvement of AE and CAS. ==> Still ongoing.
SCS
- Loop 7 is running since this morning. More loops needed.
- Missing crates for Loop 1 and 4 are finishing to be tested and will be installed back soon (Loop 1 crate already today).
- It is unclear whether the homing capability has been implemented already and can be used. AE to check.
MID
- Loop 1 still not deployed and needed. Plan is to commission the MID Opt hutch until the end of next week.
HED
- Work in tunnel ongoing
- CRL to be tested today
- HRMono- ready to start grouting.
DET
- FXE/SPB GPUs crash issue - to be checked with ITDM
XRD
- MCP in SASE2 have problems and the section has to be vented. The detector will be replaced on 23rd of Sep.
ITDM
- Office network switches will be upgraded. Intervention cancelled! Will be probably done in the Winter shutdown.
- DAQ to be upgrade to the user friendly version also in SASE3. ITDM to ask FXE whether they also want it.
|